OpenStudy (anonymous):

Solve the open sentence (Medal for best/correct answer!!) Question in comments below (I can't type it here)

OpenStudy (anonymous):

Question?

OpenStudy (anonymous):

\[-2\le n +4\le 7\]

OpenStudy (dayakar):

add -4 \[-4-2\le n+4-4 \le 7-4\]

OpenStudy (dayakar):

can u simplify it

OpenStudy (dayakar):

do u understand

OpenStudy (anonymous):

Not really, Can you break it down a bit more?

OpenStudy (janu16):

\[-6\le n \le3\] final answer

OpenStudy (janu16):

hope that helps

OpenStudy (anonymous):

Thank you :)